OpenStudy (anonymous):

Which of the following best describes the technique used to graph the equation, y = -2x + 8, using slope and y-intercept? a)Mark a point at 8 on the y-axis, move upward 2 and right 1, mark the point and draw the line. b) Mark a point at -2 on the y-axis, move downward 8 and right 1, mark the point and draw the line. c) Mark a point at -2 on the y-axis, move upward 8 and right 1, mark the point and draw the line. d) Mark a point at 8 on the y-axis, move downward 2 and right 1, mark the point and draw the line.

OpenStudy (anonymous):

well for the moving up and so forth you would use the y-int and x-int as a basis.

OpenStudy (anonymous):

What? I mean, I'm thinking it's C.

OpenStudy (johnweldon1993):

the y-intercept is...8...so you make a point at 8 first....then what?

OpenStudy (anonymous):

you are right, it is C but do you know how we can test it?

OpenStudy (johnweldon1993):

Nooooo no it's not C

OpenStudy (anonymous):

are you sure?

OpenStudy (johnweldon1993):

if it were C the y-intercept would be -2 and the equation will be y = 8x + -2

OpenStudy (johnweldon1993):

remember the slope intercept form is y=mx + b m is the slope...and b is the y-intercept

OpenStudy (anonymous):

u are right, I forgot about the the y-axis part where you marked it stays there.

OpenStudy (johnweldon1993):

right...so okay @sarah.radke you have an equation y = -2x + 8 it looks just like y = mx + b right? in that equation m is slope...and b is y-intercept so what are YOUR slope and y-intercept?

OpenStudy (anonymous):

also remember it's a negative slope so you always want it to looking from High to Low, sloping downwards.

OpenStudy (anonymous):

Oh! So -2 is the slope and the y intercept is 8?

OpenStudy (johnweldon1993):

 yes correct....and remember what your slopes look like

OpenStudy (johnweldon1993):

so now what would you do to graph your equation? what would you find and plot first?

OpenStudy (anonymous):

Oohh, okay ^.^ So it's D right? Downward two and mark 8 on the y axis?

OpenStudy (johnweldon1993):

yes you are correct :)

OpenStudy (anonymous):

Yuuussss!!! Haha thankyou so much! Gah omg I love you!

OpenStudy (johnweldon1993):

haha anytime :) Feel free to fa......nvm haha you just did fan me haha....i'll be glad to help with anything you need :)

OpenStudy (anonymous):

Haha alrightyy, thanks again!!

OpenStudy (johnweldon1993):

no problem :)